# arxiv_2024

### v1

@@ 4 Simulator performance
The following Sections describe the simulator speed. Section 4.1 first shows the raw simulator speed and peak goodput. Section 4.2 then investigates the impact on reinforcement learning workflows by evaluating the time it takes to train reinforcement learning agents through Independent PPO (IPPO) [yu2022surprising], a widely used multi-agent learning algorithm.
@@ 4.1 Simulation speed
where $A_{k}$ is the number of agents in the $k^{th}$ world, $S$ is the number of steps taken, and $\Delta T$ is the number of seconds elapsed. Figure 2 examines the scaling of the simulator as the number of simulated worlds, which represents the amount of parallelism, increases. To measure performance, we sample random batches of scenarios of size equal to the number of worlds, so that every world is a unique scenario with $K$ agents. On the left-hand side of Figure 2, we compare the performance of GPUDrive to the original Nocturne version [vinitsky2022nocturne] (CPU, no parallelism) and a CPU-accelerated version of Nocturne via Pufferlib (16 CPU cores) [suarez2024pufferlib]. Empirically, the maximum achievable AFPS of Nocturne is 15,000 (blue dotted line) though we caution that additional speedups may be possible. In contrast, GPUDrive can reach over a million ASPS on a consumer-grade GPU at 512 worlds (average agents per scenario is 60). This performance also surpasses that of Waymax [gulino2024waymax], a JAX-based simulator, where we could not run more than 32 environments in parallel due to Out of Memory (OOM) issues. Note that GPUDrive exhibits near-linear scaling of ASPS between $32$ and $128$ worlds on a datacenter-grade NVIDIA GPU and between $32$ and $256$ worlds on a consumer-grade GPU.
@@
Figure 2: Peak goodput of GPUDrive on a consumer-grade and datacenter-class GPU compared to original, CPU-based, implementations. Left: The total number of controllable agent steps per second (CASPS) as we increase the number of worlds (parallelism). Right: The total number of agent steps per second (ASPS) is the number of objects that our system computes observations for at each time step.
To ensure a fair comparison, Figure 2 (LHS) resembles the conditions used in [gulino2024waymax], where all cars, bicyclists, and pedestrians are considered as valid experience-generating agents. However, by default, our system only considers something an agent if it is necessary to move to achieve the goal. This means that cars persistently parked throughout the episode are not considered agents, which aligns correctly with what should constitute an agent. This significantly alters the characteristics of the agent distribution, leading to the right-hand side of Figure 2, where we plot the Controlled Agent Steps Per Second (CASPS). Under this definition of agents, CASPS represents the expected performance of our system when using the Waymo Open Motion Dataset (this mostly excludes parked vehicles.
@@ 4.2 End-to-end speed and performance
Figure 3 contrasts the number of steps (experience) and the corresponding time required to solve 10 scenarios from the WOMD between Nocturne and GPUDrive. For benchmarking purposes, we say a scene is solved when 95% of agents across all 10 worlds can navigate to their designated target position without colliding or going off-road across all worlds. Ceteris paribus (Details in Appendix D), GPUDrive achieves a 25-40x training speedup, solving 10 scenarios in less than 15 minutes compared to approximately 10 hours in Nocturne.
@@
As shown above, GPUDrive allows us to solve scenes in minutes. Next, we investigate how the individual scene completion time, the time it takes to solve a single scenario, changes as we increase the total number of scenarios we train in. In practice, it may be desirable to train agents on thousands of scenarios. Therefore, we ask whether it is feasible to fully leverage the simulator’s capabilities with a single GPU.
@@ 5 Conclusion
In this work, we present GPUDrive, a GPU-accelerated, data-driven simulator. GPUDrive is intended to help generate the billions of samples that are likely needed to achieve effective reinforcement learning for training multi-agent driving planners. By building atop the Madrona Engine [shacklett2023extensible], we can scale GPUDrive to hundreds of worlds containing potentially hundreds of agents leading to throughput of millions of steps per second. This throughput occurs while synthesizing complex observations such as LIDAR. We show that this throughput has consequent implications for training reinforcement learning agents, leading to the ability to train agents to solve any particular scene in minutes and in seconds when amortized across many scenes. We release the simulator and integrated trained agents to enable further research.
This paper is a first step in scaling up reinforcement learning for multi-agent planning in safety-critical, mixed human-autonomous settings. However, several important challenges remain for future work. Firstly, we have not yet identified the optimal hyperparameter settings to effectively utilize the collected data, resulting in training being the bottleneck instead of data collection time. Secondly, while the simulator is fast, collecting data for reinforcement learning leads to frequent reset calls that significantly impact throughput. Lastly, training fully human-level drivers in the simulator to navigate without crashing in any scenario remains an ongoing challenge.
Acknowledgments and Disclosure of Funding
This work is funded by the C2SMARTER Center through a grant from the U.S. DOT’s University Transportation Center Program. The contents of this report reflect the views of the authors, who are responsible for the facts and the accuracy of the information presented herein. The U.S. Government assumes no liability for the contents or use thereof. This work was also supported in part through the NYU IT High-Performance Computing resources, services, and staff expertise.
[…]
@@ A.1 Code Reproducibility
All code required to reproduce the paper is open-sourced at https://github.com/Emerge-Lab/gpudrive under release number v0.1
@@ A.2 Computational Resources
All RL experiments in this paper were run on an NVIDIA RTX 8000 or A100. Total resources for the paper correspond to less than 24 GPU-days.
Appendix B Vehicle Model
Our vehicles are driven by a kinematic bicycle model [rajamani2011vehicle] which uses the center of gravity as reference point. The dynamics are as follows. Here $(x_{t},y_{t})$ stands for the coordinate of the vehicle’s position at time $t$, $\theta_{t}$ stands for the vehicle’s heading at time $t$, $v_{t}$ stands for the vehicle’s speed at time $t$, $a$ stands for the vehicle’s acceleration and $\delta$ stands for the vehicle’s steering angle. $L$ is the distance from the front axle to the rear axle (in this case, just the length of the car) and $l_{r}$ is the distance from the center of gravity to the rear axle. Here we assume $l_{r}=0.5L$.
@@ Appendix C License Details and Accessibility
Our code is released under an MIT License. The Waymo Motion dataset is released under a Apache License 2.0. The code is available at https://github.com/Emerge-Lab/gpudrive and at release commit TODO.
@@ D.1 End-to-end performance
Figure 5: Key performance metrics as a function of training time grouped by the number of unique scenes in a batch reported in Figure 4. Left: The aggregate percentage of agents that achieved their goal. Center: The aggregate percentage of agents that collided with another vehicle. Right: The aggregate number of vehicles that crossed a road edge.

### v2

@@ 4 Simulator performance
The following Sections describe the simulator speed. Section 4.1 first shows the raw simulator speed and peak goodput (throughput achieved by the number valid number of agents in a scene). Section 4.2 then investigates the impact on reinforcement learning workflows by evaluating the time it takes to train reinforcement learning agents through Independent PPO (IPPO) (Yu et al., 2022), a widely used multi-agent learning algorithm. IPPO is implemented atop Stable Baselines 3 (Hill et al., 2018).
@@ 4.1 Simulation speed
where $A_{k}$ is the set of agents in the $k^{th}$ world, $S$ is the number of steps taken, and $\Delta T$ is the number of seconds elapsed. Figure 3 examines the scaling of the simulator as the number of simulated worlds, which represents the amount of parallelism, increases. To measure performance, we sample random batches of scenarios of size equal to the number of worlds, so that every world is a unique scenario with $K$ agents. On the left-hand side of Figure 3, we compare the performance of GPUDrive to Nocturne (Vinitsky et al., 2022) (CPU, no parallelism), a CPU-accelerated version of Nocturne via Pufferlib (16 CPU cores) (Suarez, 2024) and Waymax Gulino et al. (2024). Empirically, the maximum achievable ASPS of Nocturne is 15,000 (blue dotted line) though we note that additional speedups may be possible. GPUDrive achieves a peak ASPS of 2.3 million steps, which is 2 to 3 orders of magnitude faster compared to Nocturne. This performance also surpasses that of Waymax (Gulino et al., 2024), a JAX-based simulator, where we could not run more than 16 environments in parallel due to Out of Memory (OOM) issues.
@@
Figure 3: Peak goodput of GPUDrive on a consumer-grade and datacenter-class GPU compared to original, CPU and GPU-based, implementations using the radial filter observation. Left: The total number of agent steps per second (ASPS) is the number of objects for which our system computes observations at each time step. To ensure a fair comparison, we align the conditions with those used in (Gulino et al., 2024), where all cars, bicyclists, and pedestrians are considered valid experience-generating agents. Center: The distribution of controllable agents across 512 scenarios in the Waymo Open Motion Dataset $\mu\approx 10.8\,(\text{red line}),\sigma\approx 9.3$). Right: The total number of controllable agent steps per second (CASPS) as we increase the number of worlds (parallelism).
In addition to the ASPS, we report another metric to indicate the number of controllable agents that are stepped per second. In the Waymo Open Motion Dataset, each scenario contains a varying number of moving agents (Examples in Figure 1). By default, our system only classifies something as a controllable agent if its movement is necessary to achieve the goal. Therefore, parked cars throughout the episode are not considered controllable agents. To illustrate what this looks like, we plot the distribution of controllable agents across a subset of scenes in Figure 3.
Considering the variance of controllable agents per scene, the right-hand side of Figure 3 depicts the Controlled Agent Steps Per Second (CASPS). CASPS reflects the expected performance of our system when utilizing the Waymo Open Motion Dataset with a randomly sampled subset of scenarios. Due to the significant variability in agents throughout the dataset, the highest CASPS is notably lower than the ASPS, at around 200,000. Note that this can be improved by strategically selecting scenes, such as filtering for dense scenes with a high number of controllable agents.
Figure 4: ASPS and CASPS comparison between Radial Filter and LiDAR observation types. The radial filter is slower than lidar due to its linear scan of nearby objects. In contrast, the LiDAR observation type is GPU-accelerated, delivering significantly enhanced performance. As demonstrated in the plots, depending on the scene selection, the LiDAR can achieve a speedup of approximately 3x over the radial filter.
Lastly, we demonstrate the speed of different observation spaces in Figure 4. We observe that using LiDAR is approximately three times faster than using the radial filter observation (Details about the supported observation spaces are found in Section 3.2).
@@ 4.2 End-to-end speed and performance
Figure 5 contrasts the number of steps (experience) and the corresponding time required to solve 10 scenarios from the WOMD between Nocturne (Vinitsky et al., 2022) and GPUDrive. We compare Nocturne to GPUDrive, as it is the closest simulator in functionality, except that it is implemented on the CPU. For benchmarking purposes, we say a scene is solved when agents can navigate to their designated target position $95\%$ of the time without colliding or going off-road. Ceteris paribus (Details in Appendix D), GPUDrive achieves a 30-40x training speedup, solving 10 scenarios in less than 15 minutes compared to approximately 10 hours in Nocturne.
@@
As shown in Fig. 5, GPUDrive allows us to solve scenes in minutes. Next, we investigate how the individual scene completion time, the time it takes to solve a single scenario, changes as we increase the total number of scenarios we train on. In practice, it may be desirable to train agents on thousands of scenarios. Therefore, we ask whether it is feasible to fully leverage the simulator’s capabilities with a single GPU.
@@
4.3 Measuring remaining unsolved tasks
The Waymo Open Motion Dataset contains a total of 103,354 traffic scenarios (Ettinger et al., 2021). In this paper, we demonstrate that an agent can achieve 95% performance on a subset of 1000 scenes after 15 hours of training. Additionally, we show that the time required to solve scenarios decreases as the dataset size increases. Our analysis of the failure rates in the current best-performing policy suggests that the goal-reaching limit is around 98 %, due to mislabeled road edges in the dataset, which render some goals unreachable because they are located beyond uncrossable roads. An important direction for future work is developing agents that achieve near-perfect performance, with a goal-reaching rate approaching 100% and a 0% collision rate while addressing the limitations posed by data inaccuracies in the benchmark.
@@ 5 Conclusion
In this work, we present GPUDrive, a GPU-accelerated, data-driven simulator. GPUDrive is intended to help generate the billions of samples that are likely needed to achieve effective reinforcement learning for multi-agent driving planners. By building atop the Madrona Engine (Shacklett et al., 2023), we can scale GPUDrive to hundreds of worlds with potentially thousands of agents leading to throughput of millions of steps per second. This throughput occurs while synthesizing complex observations such as LiDAR. We show that this throughput has consequent implications for training reinforcement learning agents, leading to the ability to train agents to solve any particular scene in minutes and in seconds when amortized across many scenes. We release the simulator and integrated trained agents to enable further research.
This paper is a first step in scaling up reinforcement learning for multi-agent planning in safety-critical, mixed human-autonomous settings. However, several important challenges remain for future work. Firstly, we have not yet identified the optimal hyperparameter settings to effectively utilize the collected data, resulting in training being the bottleneck instead of data collection time. Secondly, while the simulator is fast, collecting data for reinforcement learning leads to frequent reset calls that significantly impact throughput. Lastly, training fully human-level drivers in the simulator to navigate without crashing in any scenario remains an unresolved challenge. We have shown that our simulator allows training of driving agents to achieve a 95% success rate in a few hours. However, this still does not match human capabilities or the safety standards we strive for. We hope GPUDrive will help the community tackle this open challenge.
@@ A.1 Code Reproducibility
All code required to reproduce the paper is open-sourced at
https://github.com/Emerge-Lab/gpudrive under release number v0.2.0.
@@ A.2 Computational Resources
All RL experiments in this paper were run on an NVIDIA RTX 8000 or A100. Total resources for the paper correspond to less than 2 GPU days.
Appendix B Dynamics Model
Agents are driven by a kinematic bicycle model (Rajamani, 2011) which uses the center of gravity as reference point. The dynamics are as follows. Here $(x_{t},y_{t})$ stands for the coordinate of the vehicle’s position at time $t$, $\theta_{t}$ stands for the vehicle’s heading at time $t$, $v_{t}$ stands for the vehicle’s speed at time $t$, $a$ stands for the vehicle’s acceleration and $\delta$ stands for the vehicle’s steering angle. $L$ is the distance from the front axle to the rear axle (in this case, just the length of the car) and $l_{r}$ is the distance from the center of gravity to the rear axle. Here we assume $l_{r}=0.5L$.
@@ Appendix C License Details and Accessibility
Our code is released under an MIT License. The Waymo Motion dataset is released under a Apache License 2.0. The code is available at <ANONYMIZED-DURING-REVIEW>.
@@ D.1 End-to-end performance
Goal achievement rate
Vehicle collisions
Off-road events
Figure 7: Key performance metrics as a function of training time, grouped by the number of unique scenes in a batch, as reported in Figure 6. Left: The aggregate percentage of agents that achieved their goal. Center: The aggregate percentage of agents that collided with another vehicle. Right: The aggregate number of vehicles that crossed a road edge.